# Euclidean lengths and the Culler-Shalen norms of slopes

### v1

@@ 1. Introduction
Throughout this paper, let $M$ denote a hyperbolic knot manifold, that is, a compact, orientable 3-manifold with a single torus boundary $\partial M$, whose interior admits a complete Riemannian metric of constant curvature $-1$ with finite volume. The main object studied in this paper is the set of slopes on $\partial M$. By a slope on $\partial M$, we mean the isotopy class of a simple closed curve on $\partial M$.
Let us start with the motivation and background of our research. An operation that glues a solid torus $V$ to $M$ along their boundaries is called a Dehn filling on $M$. The slope $r$ on $\partial M$, whose representative is identified with the meridian of $V$, determines the homeomorphism type of the resulting manifold. Thus, we denote the resulting manifold by $M(r)$, and say that the manifold is obtained from $M$ by Dehn filling along the slope $r$. Thurston’s well-known hyperbolic Dehn surgery theorem [24, Theorem 5.8.2] states that all but finitely many Dehn fillings on $M$ produce closed hyperbolic 3-manifolds. Accordingly, a Dehn filling on $M$ that yields a non-hyperbolic 3-manifold is called exceptional. Many works have been devoted to studying when, how many, and what types of exceptional Dehn fillings can occur.
[…]
Here we give very rough definitions of the two functions. The Euclidean length of a slope on a horotorus $T$ is defined as the minimal length of its representatives on $T$. A horotorus $T$ in $M$ appears as the image of equivariant horospheres in the hyperbolic $3$-space $\mathbb{H}^{3}$ under the universal covering map. This $T$ is naturally identified with $\partial M$ and is endowed with a Euclidean metric induced from the hyperbolic metric on $M$. Thus, the Euclidean length of curves on $\partial M$ can be defined on $T$, and the Euclidean length of a slope on $\partial M$ is defined accordingly. On the other hand, the Culler-Shalen norm is defined by using the set of characters of $SL(2,\mathbb{C})$-representations of the fundamental group $\pi_{1}(M)$. This set is called the character variety $X(\pi_{1}(M))$, which is actually regarded as an affine algebraic set. For a slope $r$, by using the trace function, a regular map $I_{\boldsymbol{r}}$ onto $\mathbb{CP}^{1}$ is defined on the smooth projective model of $X(\pi_{1}(M))$. Then the Culler-Shalen norm $\|r\|$ is defined as twice the degree of $I_{\boldsymbol{r}}$. See the next section for their precise definitions.
@@
  • If $r$ is not a strict boundary slope and $\pi_{1}(M(r))$ is cyclic, then $\|r\|$ takes the least value among non-trivial slopes [6].
  • If $r$ is not a strict boundary slope and $\pi_{1}(M(r))$ is finite, then $\|r\|$ is at most $3$ times the least value among non-trivial slopes [4].
  • If $M(r)$ does not admit a negatively-curved metric, then $\mathrm{length}_{T}(r)$ is at most $2\pi$ [3].
  • If $M(r)$ is reducible or $\pi_{1}(M(r))$ is not word-hyperbolic, then $\mathrm{length}_{T}(r)$ is at most $6$ [2, 16].
@@
In this paper, motivated by the question above, we give some inequalities between them. We first give a lower bound on the Culler-Shalen norm in terms of the Euclidean length on a horotorus when $M$ is a certain knot exterior in the $3$-sphere $S^{3}$.
@@
Suppose that $M$ is the exterior of a hyperbolic two-bridge knot or the exterior of a $(-2,3,n)$-pretzel knot with $n$ odd and at least $7$ in $S^{3}$. Then
@@
holds for any slope $r$ on $\partial M$ and for any horotorus $T$.
Let $K$ be a knot in $S^{3}$. Then $K$ is called a two-bridge knot if its bridge index (the minimum number of local maxima (or local minima) up to ambient isotopy) is two. Since two-bridge knots are alternating, a natural consequence of Menasco’s work in [18] is that a two-bridge knot is hyperbolic unless it is a $(2,p)$-torus knot. A knot $K$ in $S^{3}$ presented by a diagram obtained by putting rational tangles of the form $1/q_{1},\dots,1/q_{n}$ in line is called a pretzel knot, denoted by $P(q_{1},\dots,q_{n})$, with integers $q_{1},\dots,q_{n}$. It is known that a pretzel knot $P(-2,3,n)$ is hyperbolic unless $n=1,3,5$.
This inequality also holds for hyperbolic knots with only two strict essential surfaces, both of which are spanning surfaces, in a general $3$-manifold. See Proposition 3.2.
Our proof depends on the existence of a pair of boundary slopes with some particular property, which we can verify for such two-bridge knots and $(-2,3,n)$-pretzel knots. In fact, it is possible that the above inequality holds in a much wider class. See Section 3 for more details.
On the other hand, in general, the converse inequality does not hold. Indeed, we can find an example of a $3$-manifold whose boundary contains a slope $r$ with arbitrarily small ratio $\mathrm{length}_{T}(r)/\|r\|$ for any horotorus $T$ (see Proposition 3.4). Taking this into account, we obtain the following weaker inequality. Here, a slope on $\partial M$ is called integral if it has distance one from the meridian, and called a boundary slope if there exists an embedded essential surface in $M$ whose boundary component represents the slope. For more details on these terminologies, see the next section.
@@
When we set a meridian-longitude system on $\partial M$, the set of slopes on $\partial M$ is identified with the set of rational numbers together with $1/0$. See [21], for example. Regarding the set of non-meridional boundary slopes, denoted by $\mathcal{B}_{M}$, as a subset of the rational numbers, let $\mathrm{Diam}(\mathcal{B}_{M})$ be the diameter of the set, i.e., the difference between the greatest and the least elements.
@@
holds.
@@ Acknowledgments
The author would like to thank Masaharu Ishikawa for helpful discussions. He also thanks to Shigeru Mizushima for his help on calculations about boundary slopes for $(-2,3,n)$-pretzel knots.
@@ 2.1. Slopes
As usual, the isotopy class of an unoriented, non-trivial simple closed curve on $\partial M$ is called a slope. When we fix a particular simple closed curve on $\partial M$, called the meridian (respectively, the longitude), we refer to the slope $m$ represented by the meridian as the meridional slope (resp. longitudinal slope). The distance between two slopes is defined as the minimal geometric intersection number of their representatives. We use $\Delta(r,r^{\prime})$ to denote the distance between slopes $r$ and $r^{\prime}$. If we take a pair of a meridian and a longitude on $\partial M$, we always assume that the distance between them is one. Such a pair is called a meridian-longitude system on $\partial M$. When a meridian–longitude system, say ${m,l}$, on $\partial M$ is fixed, the set of slopes is identified with the set of rational numbers together with $1/0$: a slope $r$ corresponds to the fraction $\Delta(r,l)/\Delta(r,m)$. This gives a natural bijection. See [21], for example. We will often abuse the terminology and identify a slope with its corresponding rational number, calling it a numerical slope when necessary. From this point of view, a slope is called integral if the corresponding numerical slope lies in $\mathbb{Z}$, i.e., an integral slope has distance one from the meridional slope. We observe that, for a pair of numerical slopes $p/q$ and $r/s$, the distance $\Delta(p/q,r/s)$ and the difference $|p/q-r/s|$ as rational numbers are related as
@@
The isotopy classes of oriented, non-trivial simple closed curves on $\partial M$ are in natural bijective correspondence with the primitive elements of $H_{1}(\partial M;\mathbb{Z})$. Here, an element of $H_{1}(\partial M;\mathbb{Z})$ is called primitive if it is not an integral multiple of any other non-trivial element. Note that two primitive elements $\boldsymbol{r},\boldsymbol{r}^{\prime}\in H_{1}(\partial M;\mathbb{Z})$ correspond to the same slope if and only if $\boldsymbol{r}=\pm\boldsymbol{r}^{\prime}$. Thus, forgetting the orientations, there is a natural two-to-one map from the set of primitive elements of $H_{1}(\partial M;\mathbb{Z})$ onto the set of slopes on $\partial M$. Since a meridian and a longitude have distance one, a meridian–longitude system with orientations gives a set of generators of $H_{1}(\partial M;\mathbb{Z})$, and also of $\pi_{1}(\partial M)$. In particular, a numerical slope $p/q$ corresponds to the elements ±$\pm(p,\boldsymbol{m}+q,\boldsymbol{l})\in H_{1}(\partial M;\mathbb{Z})$, where $\boldsymbol{m}$ and $\boldsymbol{l}$ denote the elements corresponding to the oriented meridian and longitude, respectively.
@@ 2.2. Length on a horotorus
When an oriented meridian–longitude system is fixed on a horotorus $T$, we define the angle between two slopes on $T$ as follows. Let $r_{1},r_{2}$ be a pair of slopes on $T$. Take Euclidean geodesics on $T$ representing $r_{1}$ and $r_{2}$. Orient them so that they are coherent with the longitude. Then the angle between $r_{1}$ and $r_{2}$ is defined as the angle between the geodesics so oriented, measured on $T$. We measure the angle from 00 to $\pi$. In the following, for a slope $r$, we denote by $\theta_{r}$ the angle between $r$ and the meridional slope, with $0<\theta_{r}\leq\pi$.
With these settings, the next lemma holds.
@@
To give a simple proof of this lemma, we here introduce a graphical view of the length and angle for slopes. This will also be used in Subsection 4.1. As stated, we identify $\partial M$ with a horotorus $T$, and have fixed a base point on $T$. Let $\widetilde{T}$ be a component of the preimage of $T$ in the universal covering space $\mathbb{H}^{3}$ of $\mathrm{Int}(M)$. This $\widetilde{T}$ is a horosphere in $\mathbb{H}^{3}$, and so it is naturally identified with the Euclidean plane. Note that $\widetilde{T}$ is also regarded as the universal covering space of $T$, and they are locally isometric. Under the covering projection, the preimage of the base point on $T$ gives a lattice on $\widetilde{T}$. We then take a base point $O$ in the lattice. A vector from $O$ to a primitive lattice point projects down to an oriented simple closed geodesic on $T$. Since such a geodesic uniquely exists in the homology class, it determines a primitive element of $H_{1}(T;\mathbb{Z})$. Together with the relation between the slopes on $T$ and $H_{1}(T;\mathbb{Z})$, we obtain a one-to-two correspondence between the slopes on $T$ and the vectors on $\widetilde{T}$ emerging at $O$ and ending at the primitive lattice points: Two such vectors on $\widetilde{T}$ correspond to the same slope if and only if their sum is the zero vector. Now we suppose that an oriented meridian–longitude system on $\partial M$ is arbitrarily chosen and fixed. Recall that such a system gives a set of generators of $H_{1}(\partial M;\mathbb{Z})$. Using the identification of $T$ and $\partial M$, we can take two oriented Euclidean geodesics of meridional and longitudinal slope on $T$ such that both run through the base point on $T$. We set a $u$$v$ coordinate on $\widetilde{T}$ so that the vectors appearing as lifts of those two oriented geodesics starting from the origin end at $(0,1)$ and $(1,0)$, respectively. For this coordinate, by linearity, the vector appearing as the lift of the oriented geodesic on $T$ with numerical slope $p/q$ (respectively $-p/q$) corresponds to the vector ending at the point $(q,p)$ (resp. $(q,-p)$) on $\widetilde{T}$. Here we give an orientation to the geodesic on $T$ so that it is coherent with that for the longitudinal slope. In this way, we have a one-to-one correspondence between the non-meridional slopes on $\partial M$ and the vectors ending at the primitive lattice points in the half-plane $\{(u,v)\mid u>0\}$ on $\widetilde{T}$. Note that the meridional slope corresponds to the vector ending at $(0,1)$ on $\widetilde{T}$. In this view, the length of a slope on $T$ is just the length of the corresponding vector, and the angle between two slopes is just the angle between the corresponding two vectors on $\widetilde{T}$.

### v2

@@ 1. Introduction
Throughout this paper, let $M$ denote a hyperbolic knot manifold, that is, a compact, orientable 3-manifold with a single torus boundary $\partial M$, whose interior admits a complete Riemannian metric of constant curvature 1-1 with finite volume. The main object studied in this paper is the set of slopes on $\partial M$. By a slope on $\partial M$, we mean the isotopy class of an unoriented simple closed curve on $\partial M$.
Let us start with the motivation and background of our research. An operation that glues a solid torus $V$ to $M$ along their boundaries is called a Dehn filling on $M$. The slope $r$ on $\partial M$, whose representative is identified with the meridian of $V$, determines the homeomorphism type of the resulting manifold. Thus, we denote the resulting manifold by $M(r)$, and say that the manifold is obtained from $M$ by Dehn filling along the slope $r$. Thurston’s well-known hyperbolic Dehn surgery theorem [26, Theorem 5.8.2] states that all but finitely many Dehn fillings on $M$ produce closed hyperbolic 3-manifolds. Accordingly, a Dehn filling on $M$ that yields a non-hyperbolic 3-manifold is called exceptional. Many works have been devoted to studying when, how many, and what types of exceptional Dehn fillings can occur.
In the study of exceptional Dehn fillings, two functions on slopes on $\partial M$, called the Euclidean length on a horotorus $T$ and the Culler-Shalen norm on $H_{1}(\partial M;\mathbb{R})$, play important roles. See [7] for example, for a survey. In this paper, we use the notations $\mathrm{length}_{T}(\cdot)$ and $\|\cdot\|$ to denote them, respectively.
Here we give very rough definitions of the two functions. The Euclidean length of a slope on a horotorus $T$ is defined as the minimal length of its representatives on $T$. A horotorus $T$ in $M$ appears as the image of equivariant horospheres in the hyperbolic $3$-space $\mathbb{H}^{3}$ under the universal covering map. This $T$ is naturally identified with $\partial M$ and is endowed with a Euclidean metric induced from the hyperbolic metric on $M$. Thus, the Euclidean length of curves on $\partial M$ can be defined on $T$, and the Euclidean length of a slope on $\partial M$ is defined accordingly. On the other hand, the Culler-Shalen norm is defined by using the set of characters of $SL(2,\mathbb{C})$-representations of the fundamental group $\pi_{1}(M)$. This set forms a complex affine algebraic set, which is called the character variety of $M$ and denoted by $X(M)$. For a slope $r$, by using the trace function, a regular map $I_{\boldsymbol{r}}$ onto $\mathbb{CP}^{1}$ is defined on the smooth projective model of a one-dimensional component of $X(M)$. When $M$ is hyperbolic, there exists a discrete faithful representation $\pi_{1}(M)\to SL(2,\mathbb{C})$. The algebraic component $X_{0}$ of $X(M)$ containing the character of the representation is known to be one-dimensional. From twice the degree of $I_{\boldsymbol{r}}$ on $X_{0}$, a norm $\|\cdot\|_{0}$ on the real vector space $H_{1}(\partial M;\mathbb{R})$ can be obtained, which is called the Culler–Shalen norm associated to $X_{0}$. The same construction works for other one-dimensional components of $X(M)$, and in general, a seminorm on $H_{1}(\partial M;\mathbb{R})$ is obtained. Then the (total) Culler-Shalen norm $\|\cdot\|$ on $H_{1}(\partial M;\mathbb{R})$ is defined as the total sum of the seminorms on the irreducible one-dimensional components of $X(M)$. See the next section for their precise definitions.
@@
  • If $r$ is not a strict boundary slope and $\pi_{1}(M(r))$ is cyclic, then $\|r\|_{0}$ takes the least value among non-trivial slopes [8, Corollary 1.1.4].
  • If $r$ is not a strict boundary slope and $\pi_{1}(M(r))$ is finite, then $\|r\|_{0}$ is at most $5$ times the least value among non-trivial slopes [6, Theorem 2.1].
  • If $M(r)$ does not admit a negatively-curved metric, then $\mathrm{length}_{T}(r)$ is at most $2\pi$ [3, Theorem 9 (The $2\pi$” Theorem)].
  • If $M(r)$ is reducible or $\pi_{1}(M(r))$ is not word-hyperbolic, then $\mathrm{length}_{T}(r)$ is at most 66 [2, 18].
See Subsection 2.1 for details of the definitions of the terms appearing above.
@@
In this paper, motivated by the question above, we give some inequalities between them. We first give a lower bound on the (total) Culler-Shalen norm in terms of the Euclidean length on a horotorus when $M$ is a certain knot exterior in the $3$-sphere $S^{3}$.
@@
Suppose that $M$ is the exterior of a hyperbolic two-bridge knot or the exterior of a (2,3,n)(-2,3,n)-pretzel knot with $n$ odd and at least $7$ in $S^{3}$. Then
@@
holds for any slope $r$ on $\partial M$ and for any horotorus $T$. In particular, if $M$ is the exterior of a hyperbolic twist knot or the exterior of a (2,3,n)(-2,3,n)-pretzel knot with $n$ odd and at least $7$ in $S^{3}$, then the same inequality holds for $\|r\|_{0}$.
A knot $K$ in $S^{3}$ is called a two-bridge knot if its bridge index (the minimum number of local maxima (or local minima) with respect to an axis) is two. Since two-bridge knots are alternating, a natural consequence of Menasco’s work in [20] is that a two-bridge knot is hyperbolic unless it is a $(2,p)$-torus knot. Any two-bridge knot is represented by the well-known Conway diagram $C(a_{1},\dots,a_{m})$ with non-negative integers $a_{i}$. A two-bridge knot represented by the diagram $C(2,n)$ is called a twist knot. See [16] for example. A knot $K$ in $S^{3}$ presented by a diagram obtained by putting rational tangles of the form $1/q_{1},\dots,1/q_{n}$ in line is called a pretzel knot, denoted by $P(q_{1},\dots,q_{n})$, with integers $q_{1},\dots,q_{n}$. It is known that a pretzel knot P(2,3,n)P(-2,3,n) is hyperbolic unless $n=1,3,5$. See [19] for example.
Our proof depends on the existence of a pair of boundary slopes with some particular property, which we can verify for such two-bridge knots and (2,3,n)(-2,3,n)-pretzel knots. In fact, it is possible that the above inequality holds in a much wider class. For example, the inequality also holds for hyperbolic knots with only two strict essential surfaces, both of which are spanning surfaces, in a general $3$-manifold (Proposition 3.2). See Section 3 for more details.
On the other hand, in general, the converse inequality does not hold. Indeed, we can find an example of a $3$-manifold whose boundary contains a slope $r$ with arbitrarily small ratio $\mathrm{length}_{T}(r)/\|r\|$ for any horotorus $T$ (Proposition 3.4). Taking this into account, we obtain the following weaker inequality. Here, a slope on $\partial M$ is called integral if it has distance one from the meridian, and called a boundary slope if there exists an embedded essential surface in $M$ whose boundary component represents the slope. For more details on these terminologies, see the next section.
@@
When we set a meridian-longitude system on $\partial M$, the set of slopes on $\partial M$ is identified with the set of rational numbers together with $1/0$. See [23], for example. Regarding the set of non-meridional boundary slopes, denoted by $\mathcal{B}_{M}$, as a subset of the set $\mathbb{Q}$ of rational numbers, let $\mathrm{Diam}(\mathcal{B}_{M})$ be the diameter of the set, i.e., the difference between the greatest and the least elements.
@@
holds, where $m$ denotes the meridional slopes on $\partial M$.
@@ Acknowledgments
The author would like to thank Masaharu Ishikawa and Thomas Mattman for helpful discussions. He also thanks to Shigeru Mizushima for his help on calculations about boundary slopes for (2,3,n)(-2,3,n)-pretzel knots.
@@ 2.1. Slopes
As usual, the isotopy class of an unoriented, non-trivial simple closed curve on $\partial M$ is called a slope. When we fix a particular simple closed curve on $\partial M$, called the meridian (respectively, the longitude), we refer to the slope $m$ represented by the meridian as the meridional slope (resp. longitudinal slope). The distance between two slopes is defined as the minimal geometric intersection number of their representatives. We use $\Delta(r,r^{\prime})$ to denote the distance between slopes $r$ and $r^{\prime}$. If we take a pair of a meridian and a longitude on $\partial M$, we always assume that the distance between them is one. Such a pair is called a meridian-longitude system on $\partial M$. When a meridian–longitude system, say m,l{m,l}, on $\partial M$ is fixed, the set of slopes is identified with the set of rational numbers together with $1/0$: a slope $r$ corresponds to the fraction $\Delta(r,l)/\Delta(r,m)$. This gives a natural bijection. See [23], for example. We will often abuse the terminology and identify a slope with its corresponding rational number, calling it a numerical slope when necessary. From this point of view, a slope is called integral if the corresponding numerical slope lies in $\mathbb{Z}$, i.e., an integral slope has distance one from the meridional slope. We observe that, for a pair of numerical slopes $p/q$ and $r/s$, the distance $\Delta(p/q,r/s)$ and the difference |p/qr/s||p/q-r/s| as rational numbers are related as
@@
The isotopy classes of oriented, non-trivial simple closed curves on $\partial M$ are in natural bijective correspondence with the primitive elements of $H_{1}(\partial M;\mathbb{Z})$. Here, an element of $H_{1}(\partial M;\mathbb{Z})$ is called primitive if it is not an integral multiple of any other non-trivial element. Note that two primitive elements $\boldsymbol{r},\boldsymbol{r}^{\prime}\in H_{1}(\partial M;\mathbb{Z})$ correspond to the same slope if and only if $\boldsymbol{r}=\pm\boldsymbol{r}^{\prime}$. Thus, forgetting the orientations, there is a natural two-to-one map from the set of primitive elements of $H_{1}(\partial M;\mathbb{Z})$ onto the set of slopes on $\partial M$. Since a meridian and a longitude have distance one, a meridian–longitude system with orientations gives a set of generators of $H_{1}(\partial M;\mathbb{Z})$, and also of $\pi_{1}(\partial M)$. In particular, a numerical slope $p/q$ corresponds to the elements $\pm(p,\boldsymbol{m}+q,\boldsymbol{l})\in H_{1}(\partial M;\mathbb{Z})$, where $\boldsymbol{m}$ and $\boldsymbol{l}$ denote the elements corresponding to the oriented meridian and longitude, respectively.
@@ 2.2. Length on a horotorus
When an oriented meridian–longitude system is fixed on a horotorus $T$, we define the angle between two slopes on $T$ as follows. Let $r_{1},r_{2}$ be a pair of slopes on $T$. Take Euclidean geodesics on $T$ representing $r_{1}$ and $r_{2}$. Orient them so that they are coherent with the longitude. Then the angle between $r_{1}$ and $r_{2}$ is defined as the angle between the geodesics so oriented, measured on $T$. We measure the angle from 0 to $\pi$. In the following, for a slope $r$, we denote by $\theta_{r}$ the angle between $r$ and the meridional slope, with $0<\theta_{r}\leq\pi$.
With these settings, the next lemma is well-known. See [2] for example.
@@
Here, for a completeness, we include a simple proof of this lemma by using a graphical view of the length and angle for slopes. This will also be used in Subsection 4.1. As stated, we identify $\partial M$ with a horotorus $T$, and have fixed a base point on $T$. Let $\widetilde{T}$ be a component of the preimage of $T$ in the universal covering space $\mathbb{H}^{3}$ of $\mathrm{Int}(M)$. This $\widetilde{T}$ is a horosphere in $\mathbb{H}^{3}$, and so it is naturally identified with the Euclidean plane. Note that $\widetilde{T}$ is also regarded as the universal covering space of $T$, and they are locally isometric. Under the covering projection, the preimage of the base point on $T$ gives a lattice on $\widetilde{T}$. We then take a base point $O$ in the lattice. A vector from $O$ to a primitive lattice point projects down to an oriented simple closed geodesic on $T$. Since such a geodesic uniquely exists in the homology class, it determines a primitive element of $H_{1}(T;\mathbb{Z})$. Together with the relation between the slopes on $T$ and $H_{1}(T;\mathbb{Z})$, we obtain a one-to-two correspondence between the slopes on $T$ and the vectors on $\widetilde{T}$ emerging at $O$ and ending at the primitive lattice points: Two such vectors on $\widetilde{T}$ correspond to the same slope if and only if their sum is the zero vector. Now we suppose that an oriented meridian–longitude system on $\partial M$ is arbitrarily chosen and fixed. Recall that such a system gives a set of generators of $H_{1}(\partial M;\mathbb{Z})$. Using the identification of $T$ and $\partial M$, we can take two oriented Euclidean geodesics of meridional and longitudinal slope on $T$ such that both run through the base point on $T$. We set a uuvv coordinate on $\widetilde{T}$ so that the vectors appearing as lifts of those two oriented geodesics starting from the origin end at $(0,1)$ and $(1,0)$, respectively. For this coordinate, by linearity, the vector appearing as the lift of the oriented geodesic on $T$ with numerical slope $p/q$ (respectively p/q-p/q) corresponds to the vector ending at the point $(q,p)$ (resp. (q,p)(q,-p)) on $\widetilde{T}$. Here we give an orientation to the geodesic on $T$ so that it is coherent with that for the longitudinal slope. In this way, we have a one-to-one correspondence between the non-meridional slopes on $\partial M$ and the vectors ending at the primitive lattice points in the half-plane $\{(u,v)\mid u>0\}$ on $\widetilde{T}$. Note that the meridional slope corresponds to the vector ending at $(0,1)$ on $\widetilde{T}$. In this view, the length of a slope on $T$ is just the length of the corresponding vector, and the angle between two slopes is just the angle between the corresponding two vectors on $\widetilde{T}$.